Project

General

Profile

Update 15 January 2016

News: Optics Analysis
Added by Kyberd, Paul 3 months ago

Dear All,

The “big” event this week was the first formal review of the Optics Analysis by an external panel of experts. This was both a success and an important step for the collaboration.

The excellence of the material, which is collected on the review page (http://micewww.pp.rl.ac.uk/projects/governance/wiki/2016-01-14-15-OpticsReview), speaks for itself. The document prepared to brief the reviewers (Status of the MICE Optics and Analysis rev 1), the presentations and the material prepared in response to the reviewers’ questions will be an important resource in the weeks to come.

I would like to congratulate Chris Rogers and the analysers not only for the excellent talks but also for demonstrating an impressive grasp of the accelerator, detector and experimental physics of MICE and of the outstanding issues that need to be addressed; their command of detail, apparently effortless, was was recognised by the reviewers.

The two days of the review were both enjoyable and inspiring; recording our achievements to date, the capabilities we have at Step IV and the potential for impact with the demonstration of ionization cooling.

A brief status report follows …

With very best wishes for 2016 …

Ken

Status update:

Magnets:
• Absorber/focus-coil module:
⁃ A leak, perhaps the leak, in the hydrogen system has been identified. It has been located in the indium seal of the downstream window. This is excellent news. The module now needs to be taken to R9 for the hydrogen vessel to be withdrawn and a repair effected. The move of the AFC from the Hall to R9 is scheduled to take place on Friday 22Jan16.
• Spectrometer solenoids:
⁃ Progress has been made on the design and specification of the QD/QP system required to allow operation of SSU and SSD at Step IV. Over the coming two weeks the system will be specified. Discussions about the implementation will then gather momentum.
⁃ The SSD repair plan is being developed. Clearly a difficult problem. A joined up approach is being taken with potential manufacturers being considered on both sides of the Atlantic in a coherent effort between US personnel and personnel at RAL.
• Decay solenoid:
⁃ The repaired FUG power supply is still expected to be returned to RAL by the end of February 2016. Once receipted, the module will need to be installed and commissioned.
⁃ The Linde fridge has been serviced. Next week, the helium compressor that serves the Linde fridge will be serviced.

MLCR:
• The controls and monitoring network re-configuration is in progress. This is necessary to separate the controls from the daq network and to increase the robustness of communications between the various critical modules on the controls network.


Comments